# Principled Instructions Are All You Need for Questioning LLaMA-1/2, GPT-3.5/4

### 26 Dec 2023 v1

@@ Abstract
This paper introduces 26 guiding principles designed to streamline the process of querying and prompting large language models. Our goal is to simplify the underlying concepts of formulating questions for various scales of large language models, examining their abilities, and enhancing user comprehension on the behaviors of different scales of large language models when feeding into different prompts. Extensive experiments are conducted on LLaMA-1/2 (7B, 13B and 70B), GPT-3.5/4 to verify the effectiveness of the proposed principles on instructions and prompts design. We hope that this work provides a better guide for researchers working on the prompting of large language models. Project page is available at https://github.com/VILA-Lab/ATLAS.
@@ 1 Introduction
We elaborate the principled instructions for LLM prompting, provide further motivation, and detail several specific designing principles in Section 3. In Section 4 we show experimentally that the proposed principles can produce higher quality, more concise, factual and less complicated or intricate responses than standard prompts for LLMs. Specifically, with the manually-designed ATLAS benchmark, which includes multiple questions for each principle, the specialized prompts we introduced have enhanced both the quality and accuracy of the LLM responses by an average of 57.7% and 67.3%, respectively, when applied to GPT-4. Furthermore, the improvements are more pronounced with the increase in model size, for example, the performance gains when moving from LLaMA-2-7B to GPT-4 exceed 40%.
@@ 2 Related Work
Prompting. Prompting, as a distinct aspect of interacting with language models and its simplicity with no need to fine-tune the model, has evolved into a nuanced field of study, highlighting the intricate relationship between user inputs and LLM responses. Early explorations, such as those by [19], delved into how varying prompt designs could dramatically influence the performance and outputs of language models, marking the birth of prompt engineering. This area rapidly expanded, uncovering the critical role of prompts in few-shot and zero-shot learning scenarios, exemplified by [2] work with GPT-3, where strategically crafted prompts enabled the model to perform tasks with minimal prior examples. Beyond mere task instruction, recent studies have shifted towards understanding the semantic and contextual nuances in prompts, examining how subtle changes can lead to significantly different responses from the LLM.
@@ 3.2 Overview
No need to be polite with LLM so there is no need to add phrases like “please”, “if you don’t mind”, “thank you”,
“I would like to”, etc., and get straight to the point.
@@
Add to your prompt the following phrase “Ensure that your answer is unbiased and does not rely on stereotypes”.
@@
Allow the model to elicit precise details and requirements from you by asking
you questions until he has enough information to provide the needed output
(for example, “From now on, I would like you to ask me questions to…”).
@@
the following phrase: “Teach me the [Any theorem/topic/rule name] and include a test at the end, but don’t
give me the answers and then tell me if I got the answer right when I respond”.
@@
only improve the user’s grammar and vocabulary and make sure it sounds natural. You should not change the
writing style, such as making a formal paragraph casual”.
@@
o I’m providing you with the beginning [song lyrics/story/paragraph/essay…]: [Insert lyrics/words/sentence]’. Finish it based on the words provided. Keep the flow consistent.
@@
o Please use the same language based on the provided paragraph[/title/text /essay/answer].
Table 1: Overview of 26 prompt principles.
@@
    o Explain to me like I’m 11 years old
    o Explain to me as if I’m a beginner in [ field ]
    o “Write the [essay/text/paragraph] using simple English like you’re explaining something to a 5-year-old”
@@
Add to your prompt the following phrase “Ensure that your answer is unbiased and does not rely on stereotypes.”
@@
    o “Please use the same language based on the provided paragraph.[/title/text /essay/answer]”
@@
    o “Teach me the [Any theorem/topic/rule name] and include a test at the end, but don’t give me the answers and then tell me if I got the answer right when I respond”
@@
    o “From now on, I would like you to ask me questions to…”.
@@
To write an essay /text /paragraph /article or any type of text that should be detailed: “Write a detailed [essay/text/paragraph] for me on [topic] in detail by adding all the information necessary”.
@@
To correct/change specific text without changing its style: “Try to revise every paragraph sent by users. You should only improve the user’s grammar and vocabulary and make sure it sounds natural. You should not change the writing style, such as making a formal paragraph casual.”
@@ 4.1 Setup and Implementation Details
All our evaluation is performed on ATLAS [18], a manually crafted benchmark for principled prompt evaluation. For each principle, it contains 20 human-selected questions with and without the principled prompts. Following [10, 25], we evaluate the various scales of LLM outputs by human evaluation.
@@
Figure 5: Correctness improvement of LLM response quality after employing the introduced principles on prompts. small-scale indicates the 7B models, medium-scale indicates the 13B models and large-scale indicates the 70B and GPT-3.5/4 models.
@@ 4.2 Models and Metrics
We use instruction finetuned LLaMA-1-{7, 13}, LLaMA-2-{7, 13}, off-the-shelf LLaMA-2-70B-chat, GPT-3.5 (ChatGPT) and GPT-4 as our base models. We group these models into different scales: small-scale (7B models), medium-scale (13B) and large-scale (70B, GPT-3.5/4). We evaluate these models in two settings: Boosting and Correctness. They are employed together to provide a comprehensive understanding of a model’s performance.
  • Boosting. We assess the enhancement in the quality of responses from different LLMs via human evaluation after applying the outlined prompt principles. The original, unmodified prompts act as a benchmark for measuring this enhancement. Demonstrating boosting confirms that a model’s performance has improved due to the use of structured, principled instructions, as shown in Fig. 2.
  • Correctness. The concept of correctness refers to the precision of the model’s outputs or responses, ensuring they are accurate, relevant, and devoid of errors. Human evaluators are utilized to gauge this aspect, which is crucial for verifying the model’s accuracy. Correctness is a testament to the model’s ability to generate outputs that align with the expected standards of accuracy, as shown in Fig. 3.
@@ 4.3 Results
Figure 7: Correctness improvement score on the ATLAS dataset.
Figure 8: Illustration of heatmap for LLMs improvement percentages.
Figure 9: Illustration of heatmap for LLMs correctness percentages.
@@ 4.3.1 Results on small, medium and large-scale LLMs
Boosting. The results of improvement after employing the introduced principles are shown in Fig. 5. Generally, all principles can bring a significant improvement on the three scales of LLMs. In the cases of principles 2, 5, 15, 16, 25 and 26, the large-scale models get the most improvement by the principled prompts.
Correctness. As shown in Fig. 5, the employment of all principles typically results in improvements exceeding 20% on the averaged various models. In particular, for small and medium scale models, the improvement can basically reach between 20% and 30%, and for large models, the improvement can reach more than 50%.
@@ 4.3.2 Results on individual LLMs
Correctness. Fig. 7 illustrates the enhancements in accuracy across different sizes of LLMs. From LLaMA-2-13B, LLaMA-2-70B-chat to GPT-3.5 and GPT-4, there is a noticeable trend: the larger the model, the greater the increase in correctness. Fig. 9 further presents the correctness enhancements by each principle.
@@ 4.3.3 More examples on various scales of LLMs
We present additional examples for both small and medium-scale LLMs, as illustrated in Fig. 10 and 11 for the small-scale LLaMA-2-7B, and Fig. 12 and 13 for the medium-scale LLaMA-2-13B. Empirically, the use of the proposed principles on prompts has demonstrably enhanced the accuracy of the responses generated by these models.
@@ 6 Limitations and Discussion
Despite our efforts in evaluating these principles on seven distinct language models, it is crucial to acknowledge that models with architectures different from those tested might respond in different ways to these principles. Additionally, our assessment of improvement and correctness percentages was based on a limited selection of questions. Expanding the question set in future research could yield more generalized findings and offer deeper insights into the applicability of each principle.

### 18 Jan 2024 v2

@@ Abstract
This paper introduces 26 guiding principles designed to streamline the process of querying and prompting large language models. Our goal is to simplify the underlying concepts of formulating questions for various scales of large language models, examining their abilities, and enhancing user comprehension on the behaviors of different scales of large language models when feeding into different prompts. Extensive experiments are conducted on LLaMA-1/2 (7B, 13B and 70B), GPT-3.5/4 to verify the effectiveness of the proposed principles on instructions and prompts design. We hope that this work can provide a better guide for researchers working on the prompting of large language models. Project page is available at https://github.com/VILA-Lab/ATLAS.
@@ 1 Introduction
We elaborate the principled instructions for LLM prompting, provide further motivation, and detail several specific designing principles in Section 3. In Section 4 we show experimentally that the proposed principles can produce higher quality, more concise, factual and less complicated or intricate responses than standard prompts for LLMs. Specifically, with the manually-designed ATLAS benchmark, which includes multiple questions for each principle, the specialized prompts we introduced have enhanced both the quality and accuracy of the LLM responses by an average of 57.7% and 36.4%, respectively, when applied to GPT-4. Furthermore, the improvements are more pronounced with the increase in model size, for example, the performance gains when moving from LLaMA-2-7B to GPT-4 exceed 20%.
@@ 2 Related Work
Prompting. Prompting [20, 12, 25, 27, 14], as a distinct aspect of interacting with LLMs and its simplicity with no need to fine-tune the model, has evolved into a nuanced field of study, highlighting the intricate relationship between user inputs and LLM responses. Early explorations, such as those by [20], delved into how varying prompt designs could dramatically influence the performance and outputs of language models, marking the birth of prompt engineering. This area rapidly expanded, uncovering the critical role of prompts in few-shot and zero-shot learning scenarios, exemplified by [2] work with GPT-3, where strategically crafted prompts enabled the model to perform tasks with minimal prior examples. Beyond mere task instruction, recent studies have shifted towards understanding the semantic and contextual nuances in prompts, examining how subtle changes can lead to significantly different responses from the LLM.
@@ 3.2 Overview
If you prefer more concise answers, no need to be polite with LLM so there is no need to add phrases like
“please”, “if you don’t mind”, “thank you”, “I would like to”, etc., and get straight to the point.
@@
Add to your prompt the following phrase “Ensure that your answer is unbiased and avoids relying on stereotypes.”
@@
Allow the model to elicit precise details and requirements from you by asking you questions until he has
enough information to provide the needed output (for example, “From now on, I would like you to ask me
questions to …”).
@@
the following phrase: “Teach me any [theorem/topic/rule name] and include a test at the end, and let me know if
my answers are correct after I respond, without providing the answers beforehand.”
@@
only improve the user’s grammar and vocabulary and make sure it sounds natural. You should maintain the original
writing style, ensuring that a formal paragraph remains formal.”
@@
o I’m providing you with the beginning [song lyrics/story/paragraph/essay…]: [Insert lyrics/words/sentence]. Finish it based on the words provided. Keep the flow consistent.
@@
o Use the same language based on the provided paragraph[/title/text /essay/answer].
Table 1: Overview of 26 randomly ordered prompt principles.
@@
    o Explain to me like I’m 11 years old.
    o Explain to me as if I’m a beginner in [ field ].
    o “Write the [essay/text/paragraph] using simple English like you’re explaining something to a 5-year-old.”
@@
Add to your prompt the following phrase “Ensure that your answer is unbiased and avoids relying on stereotypes.”
@@
    o “Use the same language based on the provided paragraph [/title/text/essay/answer].”
@@
    o “Teach me the [Any theorem/topic/rule name] and include a test at the end, and let me know if my answers are correct after I respond, without providing the answers beforehand.”
@@
    o “From now on, I would like you to ask me questions to …”
@@
To write an essay /text /paragraph /article or any type of text that should be detailed: “Write a detailed [essay/text/paragraph] for me on [topic] in detail by adding all the necessary information.”
@@
To correct/change specific text without changing its style: “Try to revise every paragraph sent by users. You should only improve the user’s grammar and vocabulary and make sure it sounds natural. You should maintain the original writing style, ensuring that a formal paragraph remains formal.”
@@ 4.1 Setup and Implementation Details
All our evaluation is performed on ATLAS [19], a manually crafted benchmark for principled prompt evaluation. It contains a standard subset featuring questions across various domains, along with a challenging subset dedicated to reasoning and other complex tasks. In our evaluation, we utilize a single response for each question. For each principle and the challenging subset, it contains 20 human-selected questions with and without the principled prompts. Similar to [10, 26], we compare each pair of responses from the same instructions with and without principles, and evaluate the various scales of LLM outputs by human evaluation.
@@
Figure 5: Absolute correctness of LLM response quality after employing the introduced principles on prompts. small-scale indicates the 7B models, medium-scale indicates the 13B models and large-scale indicates the 70B and GPT-3.5/4 models.
Figure 6: Relative correctness improvement of LLM response quality after employing the introduced principles on prompts. small-scale indicates the 7B models, medium-scale indicates the 13B models and large-scale indicates the 70B and GPT-3.5/4 models.
@@ 4.2 Models and Metrics
We use instruction finetuned LLaMA-1-{7, 13}, LLaMA-2-{7, 13}, off-the-shelf LLaMA-2-70B-chat, GPT-3.5 (ChatGPT) and GPT-4 as our base models. We group these models into different scales: small-scale (7B models), medium-scale (13B) and large-scale (70B, GPT-3.5/4). We evaluate these models in two settings: Boosting and Correctness. They are employed together to provide a comprehensive understanding of a model’s performance. For correctness, we specifically utilize complex reasoning tasks to accurately gauge the precision of the models’ outputs, contrasting with our evaluation for boosting, where simpler tasks are employed to effectively measure quality improvements. This distinction ensures a better reflection of the true capabilities for different scales of models and the effect of the principles for prompts. Since we use questions that typically involve complex reasoning tasks for correctness, some principles are not applicable including principles 14, 15, 21, 22, 23. For instance, “Suppose $a$ and $b$ are positive real numbers with $a>b$ and $ab=8$. Find the minimum value of $\frac{a^{2}+b^{2}}{a-b}$.
  • Boosting. The result of boosting refers to the percentage increase in response quality across a set of questions when the proposed principles are applied. We assess the enhancement in the quality of responses from different LLMs via human evaluation after applying the outlined prompt principles. The original, unmodified prompts act as a baseline for measuring this enhancement. Demonstrating boosting confirms that a model’s performance has improved due to the use of structured, principled instructions, as shown in Fig. 2.
  • Correctness. The concept of correctness refers to the precision of the model’s outputs or responses, ensuring they are accurate, relevant, and devoid of errors. We consider both absolute and relative correctness accuracy. Human evaluators are utilized to gauge this aspect, which is crucial for verifying the model’s accuracy. Correctness is a testament to the model’s ability to generate outputs that align with the expected standards of accuracy, as shown in Fig. 3.
@@ 4.3 Results
Figure 7: Boosting score across various LLMs on the ATLAS dataset.
Figure 8: Absolute correctness score on the ATLAS dataset.
Figure 9: Relative correctness improvement score on the ATLAS dataset.
Figure 10: Illustration of heatmap for LLMs boosting percentages.
Figure 11: Illustration of heatmap for absolute correctness percentages.
Figure 12: Illustration of heatmap for relative correctness improvement percentages.
@@ 4.3.1 Results on small, medium and large-scale LLMs
Boosting. The results of improvement after employing the introduced principles are shown in Fig. 5. Generally, all principles can bring a significant improvement on the three scales of LLMs. In the cases of principles 2, 5, 15, 16, 25 and 26, the large-scale models get the most improvement by the principled prompts. Particularly, for principle 14, as shown in Fig. 5, it has improved all questions it is applied to.
Correctness. (1) Absolute accuracy: we examine the absolute performance when employing the principles on various scales of models. Generally, these models achieve 20$\sim$40accuracy on the averaged performance, as shown in Fig. 5. In particular, for small and medium scale models, the accuracy can basically reach between 10% and 40%, and for large models, the accuracy can reach more than 40%. (2) Relative accuracy: Fig. 6 illustrates that applying the principles generally leads to a performance increase of over 10% across different models on average. For larger models, this enhancement can surpass 20%.
@@ 4.3.2 Results on individual LLMs
Correctness. Fig. 8 illustrates the absolute correctness accuracy and Fig. 9 shows the relative enhancements in accuracy across different sizes of LLMs. From LLaMA-2-13B, LLaMA-2-70B-chat to GPT-3.5 and GPT-4, there is a noticeable trend: the larger the model, the greater the increase in correctness improvement. Fig. 11 and Fig. 12 further present the absolute and relative correctness enhancements by each principle.
@@ 4.3.3 More examples on various scales of LLMs
We present additional examples for both small and medium-scale LLMs, as illustrated in Fig. 13 and 14 for the small-scale LLaMA-2-7B, and Fig. 15 and 16 for the medium-scale LLaMA-2-13B. Empirically, the use of the proposed principles on prompts has demonstrably enhanced the accuracy of the responses generated by these models.
@@ 6 Limitations and Discussion
Despite our efforts in evaluating these principles on seven distinct language models, it is crucial to acknowledge that models with architectures different from those tested might respond in different ways to these principles. Additionally, our assessment of improvement and correctness percentages was based on a limited selection of questions. Expanding the question set in future research could yield more generalized findings and offer deeper insights into the applicability of each principle. Furthermore, the criteria and results may vary across various personnel assessments on the model responses.